A bicycle is traveling at 20 miles per hour.
How many feet will it cover in 30 seconds?
Round your answer to the nearest tenth of a foot.

Found 2 solutions by math_tutor2020, Theo:
Answer by math_tutor2020(3838)


1 mile = 5280 ft
1 hour = 60 min
1 min = 60 sec

Converting from mph to ft per sec


"miles" units cancel

"hours" units cancel


"minutes" units cancel
The only thing left will be "ft" up top and "sec" down below
That gives the speed unit ft/sec aka "feet per second", often abbreviated to fps.

Let's finish up the calculation






which is approximate; the '3's go on forever.

The person biking at 20 mph is moving at roughly 29.33333333 ft/sec

distance = rate*time
distance = (29.33333333 ft/sec)*(30 sec)
distance = 879.9999999
distance = 880 feet is the final answer.

------------------------------------------------

Another approach:

1 hour = 60 min
1 min = 60 sec
Those two equations tell us that
1 hour = 60*60 = 3600 sec

Meaning that 1 second is 1/3600 of an hour.
30 seconds is 30*(1/3600) = 30/3600 = 1/120 of an hour.

distance = rate*time
distance = (20 mph)*(1/120 of an hour)
distance = 20/120 miles
distance = 1/6 of a mile.


Then,
1 mile = 5280 ft
(1/6)*1 mile = (1/6)*5280 ft
1/6 of a mile = 880 feet


------------------------------------------------


Answer: 880 feet

Answer by Theo(13342)

the bicycle is traveling at 20 miles per hour.
20 miles = 20 * 5280 = 105600 feet.
1 hour = 3600 seconds.
the bicycle is therefore traveling 105600 in 3600 seconds.
that's equal to 105600/3600 = 29.333333333... feet per second.
therefore, the bicycle travels 30 * 29.333333333... = 880 feet in 30 seconds.
that's the number of feet the bicycle traveled in 30 seconds.